OpenStudy (anonymous):

Media Services charges $20 for a phone and $30 month for its economy plan. ?The equation c = 30t + 20 describes the total cost c of operating a Media Services phone for t months The total cost for 7 months of service? If a customer has only $90 available how many months of service can she receive?

OpenStudy (anonymous):

c=30t+20 C=30(7)+20 C=210 + 20 C=230

OpenStudy (anonymous):

C=90 90=30t+20 70=30t t=2 only two months will be available and 10 dollars would be left over

OpenStudy (anonymous):

That's $50 less per 7 months than my Metro PCS.. !!